# Visual Highlighting for Situated Brushing and Linking

### v1

@@ 6.1 Observations and lessons learned
Since Arrow clearly has the lowest results for general experience and enjoyment, we can declare it the least preferred technique for GQ2 and GQ3, at least, for our scenario. This observation is also reflected in the comments of our participants, who described Arrow as "irritating" (P1), "annoying" (P7) or even "exhausting for the eyes" (P13). Gruenefeld et al. [$*{}^{*}$18] had similar task performance results, but found a lower workload for Arrow, which could be due to our scenario with its many targets. Our animated Arrow resulted in worse results than the static arrows of Wieland et al. [WGRF22] and Yu et al. [$*{}^{*}$20], which could be due to its usage as a baseline condition based on familiarity reasons. Since Wieland et al. [WGRF22] only considered a single target at a time, we assume that their arrow approach would also result in clutter for multiple targets. One reason might be the clutter introduced by Arrow that affects its ability to highlight many targets. P12 observed that "[Arrow is] obstructing too much of the view, especially when there are multiple targets" (P12). However, Link was not perceived as cluttered, suggesting that animation might causes poor scalability. Additionally, all previous approaches [$*{}^{*}$18, WGRF22, $*{}^{*}$20] included estimation tasks that could have led to better results for Arrow if applied in our context. One could modify Arrow such that it is not affected by head movements if many targets are present. However, the clutter of Arrow remains a major drawback.
@@ 6.3 Future directions
We intentionally used baseline highlighting techniques established in brushing and linking and attention guidance. Based on our results, there are clear avenues for improving each of the four techniques. Although we did not directly evaluate the matter, Color lacks a way to perceive the actual target (since it is distorted by the color). A modified Outline with increased contrast and outline width could inherit most of the benefits of Color, while not (fully) occluding the underlying targets. For Link, we suggest that adjusting size based on distance, as well as adding a way of perceiving lines behind the tablet or user, would further increase the already encouraging results. However, care has to be taken to avoid excessive visual clutter [$*{}^{*}$19]. We assume that Arrow might be more beneficial if the user’s movements did not have an impact on its path, or if used as a supplementary indicator combined with Color or Outline. In general, it seems likely that a hybrid technique can overcome the weaknesses of individual techniques, provided that we can avoid too much clutter. However, the optimal choice will likely depend on the scenario. For example, applying Color to large objects can be visually overwhelming, even if Color was the most favored overall. Future work may seek to devise an adaptive highlighting technique to tune the visual characteristics of the highlighting to the current viewing situation.
Future work should also investigate brushing. If brushing in reverse, i.e., brushing physical referents, an analyst could compare the values of nearby referent after physically navigating to an interesting real-world area, such as inspecting the bargain bin in the supermarket. Several techniques could support brushing in this manner, including spatial filtering [FMS93, $*{}^{*}$00], aggregation [$*{}^{*}$16], and selection [$*{}^{*}$20] techniques. We leave the exploration of this inverse brushing for future work.
@@ 6.4 Limitations
Another limitation is our use of VR to simulate AR, even if transferability from VR to AR has been demonstrated for other tasks [LBHB09, $*{}^{*}$13]. We assume that any differences observed between the highlighting techniques in our study will still translate into AR, and the reasons for using one technique over the other remain the same. However, it is likely that their effectiveness in AR will largely be dictated by the technical capabilities of the AR headset and tracking solution [CFSS23]. For example, tracking drift may cause noticeable misalignment in object silhouettes when using Color, which may justify the use of other techniques less sensitive to this issue like Link or Arrow. The study of these considerations lies firmly outside the scope of this work.

### v3

@@ 6.1 Observations and lessons learned
Since Arrow clearly has the lowest results for general experience and enjoyment, we can declare it the least preferred technique for GQ2 and GQ3, at least, for our scenario. This observation is also reflected in the comments of our participants, who described Arrow as "irritating" (P1), "annoying" (P7) or even "exhausting for the eyes" (P13). Gruenefeld et al. [GLH18] had similar task performance results, but found a lower workload for Arrow, which could be due to our scenario with its many targets. Our animated Arrow resulted in worse results than the static arrows of Wieland et al. [WGRF22] and Yu et al. [YLF20], which could be due to its usage as a baseline condition based on familiarity reasons. Since Wieland et al. [WGRF22] only considered a single target at a time, we assume that their arrow approach would also result in clutter for multiple targets. One reason might be the clutter introduced by Arrow that affects its ability to highlight many targets. P12 observed that "[Arrow is] obstructing too much of the view, especially when there are multiple targets" (P12). However, Link was not perceived as cluttered, suggesting that animation might causes poor scalability. Additionally, all previous approaches [GLH18, WGRF22, YLF20] included estimation tasks that could have led to better results for Arrow if applied in our context. One could modify Arrow such that it is not affected by head movements if many targets are present. However, the clutter of Arrow remains a major drawback.
@@ 6.3 Future directions
We intentionally used baseline highlighting techniques established in brushing and linking and attention guidance. Based on our results, there are clear avenues for improving each of the four techniques. Although we did not directly evaluate the matter, Color lacks a way to perceive the actual target (since it is distorted by the color). A modified Outline with increased contrast and outline width could inherit most of the benefits of Color, while not (fully) occluding the underlying targets. For Link, we suggest that adjusting size based on distance, as well as adding a way of perceiving lines behind the tablet or user, would further increase the already encouraging results. However, care has to be taken to avoid excessive visual clutter [HRD19]. We assume that Arrow might be more beneficial if the user’s movements did not have an impact on its path, or if used as a supplementary indicator combined with Color or Outline. In general, it seems likely that a hybrid technique can overcome the weaknesses of individual techniques, provided that we can avoid too much clutter. However, the optimal choice will likely depend on the scenario. For example, applying Color to large objects can be visually overwhelming, even if Color was the most favored overall. Future work may seek to devise an adaptive highlighting technique to tune the visual characteristics of the highlighting to the current viewing situation.
Future work should also investigate brushing. If brushing in reverse, i.e., brushing physical referents, an analyst could compare the values of nearby referent after physically navigating to an interesting real-world area, such as inspecting the bargain bin in the supermarket. Several techniques could support brushing in this manner, including spatial filtering [FMS93, JLB00], aggregation [TOK16], and selection [SCZ20] techniques. We leave the exploration of this inverse brushing for future work.
@@ 6.4 Limitations
Another limitation is our use of VR to simulate AR, even if transferability from VR to AR has been demonstrated for other tasks [LBHB09, LRM13]. We assume that any differences observed between the highlighting techniques in our study will still translate into AR, and the reasons for using one technique over the other remain the same. However, it is likely that their effectiveness in AR will largely be dictated by the technical capabilities of the AR headset and tracking solution [CFSS23]. For example, tracking drift may cause noticeable misalignment in object silhouettes when using Color, which may justify the use of other techniques less sensitive to this issue like Link or Arrow. The study of these considerations lies firmly outside the scope of this work.